13-)The 2011 values for a pension fund are provided in the table below, in millions of dollars.

                    Pension Assets              

Beginning value on 1 January

Change in value in 2011

$50.11

6.30%

Pension Liabilities        

Beginning value on 1 January

Change in value in 2011   

$47.23

30%

Multiple choice question: Which of the following is closest to the surplus return for this fund?

      a)     -28.1%

  1. ) -2.1%
  2. )   7.3%
  3. ) -6.8%

-- PLEASE SHOW YOUR WORK/
Answer #1

There is something wrong in question, as per my findings, change in value of Pension liabilities should have been given 8.4% and i have answered according to that.

1. Beginning Surplus = $50.11-47.23 = $2.88
Pension assets at end of 2011 = $50.11 x 1.063 = $53.27
Pension Liabilities at end of 2011 = $47.23 x 1.084 = $51.20
Ending Surplus = $53.27-51.20 = $2.07

Change in Surplus = (2.88-2.07)/2.88 = -28.1%

Answer is a) -28.1%